OpenStudy (jamaljunior):

Two planes left the same airport at the same time traveling in opposite directions. The first plane traveled at an average rate of 30 mph slower than the second plane. After 3 hours, the planes were 3150 miles apart. Enter an equation that can be used to determine when the planes are 3150 miles apart, where x represents the average rate of the second plane. Enter the average rate of change of the slower plane

OpenStudy (jamaljunior):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get with this?

OpenStudy (jamaljunior):

no where

OpenStudy (jamaljunior):

idk where to start

jimthompson5910 (jim_thompson5910):

which plane is traveling faster? the first or second?

OpenStudy (jamaljunior):

second

jimthompson5910 (jim_thompson5910):

`The first plane traveled at an average rate of 30 mph slower than the second plane` `x represents the average rate of the second plane` speed of first plane = x-30 speed of second plane = x do you agree?

OpenStudy (jamaljunior):

yes

jimthompson5910 (jim_thompson5910):

if the speed of the second plane is x and it travels for 3 hours, then what distance does it travel in terms of x?

jimthompson5910 (jim_thompson5910):

hint: distance = speed*time

OpenStudy (jamaljunior):

x+30?

jimthompson5910 (jim_thompson5910):

second plane distance = unknown speed = x time = 3 distance = speed*time distance = x*3 distance = 3x make sense?

OpenStudy (jamaljunior):

yes

jimthompson5910 (jim_thompson5910):

what is the distance for the first plane?

OpenStudy (jamaljunior):

x-30

jimthompson5910 (jim_thompson5910):

that's the speed

jimthompson5910 (jim_thompson5910):

it travels for 3 hours what is the distance?

OpenStudy (jamaljunior):

-30 x 3 = -90

jimthompson5910 (jim_thompson5910):

first plane distance = unknown speed = x-30 time = 3 distance = speed*time distance = (x-30)*3 distance = 3x-90 make sense?

OpenStudy (jamaljunior):

yes

jimthompson5910 (jim_thompson5910):

(distance of first plane) + (distance of second plane) = total distance (3x) + (3x - 90) = 3150 3x + 3x - 90 = 3150 6x - 90 = 3150 I'll let you solve for x

OpenStudy (jamaljunior):

39.93

OpenStudy (jamaljunior):

6x - 90 = 3150 -6 -6 ---------------- x - 90 = 3144 90/90 3144/90 (39.93)

OpenStudy (jamaljunior):

is that it?

jimthompson5910 (jim_thompson5910):

6x - 90 = 3150 6x - 90+90 = 3150+90 .... add 90 to both sides 6x = 3240 x = ???

OpenStudy (jamaljunior):

540

jimthompson5910 (jim_thompson5910):

yes x = 540

jimthompson5910 (jim_thompson5910):

so x-30 = 540-30 = 510

OpenStudy (jamaljunior):

how do i put it into an equation?

jimthompson5910 (jim_thompson5910):

speed of first plane = 510 mph speed of second plane = 540 mph

OpenStudy (jamaljunior):

so the average rate of change is 30mph?

jimthompson5910 (jim_thompson5910):

speed = average rate of change

OpenStudy (jamaljunior):

oh okay

OpenStudy (jamaljunior):

how do i put it into an equation?

jimthompson5910 (jim_thompson5910):

we just solved the equation above

OpenStudy (jamaljunior):

yes but i have to answer it in an equation

OpenStudy (jamaljunior):

Enter an equation that can be used to determine when the planes are 3150 miles apart, where x represents the average rate of the second plane.

jimthompson5910 (jim_thompson5910):

the equation is shown above

OpenStudy (jamaljunior):

this? x-30 = 540-30 = 510

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

the first equation I wrote out

OpenStudy (jamaljunior):

x = 540

jimthompson5910 (jim_thompson5910):

scroll up more

OpenStudy (jamaljunior):

6x - 90 = 3150 6x - 90+90 = 3150+90 .... add 90 to both sides 6x = 3240 x = ???

jimthompson5910 (jim_thompson5910):

they're probably looking for 6x-90 = 3150 or 3x+3x-90 = 3150

OpenStudy (jamaljunior):

thats my guess. ok thanks again jim

jimthompson5910 (jim_thompson5910):

no problem